Ali will be 6y years old in 6 years time.How old was Ali 4y years ago?
A.6y-6
B.10y-6
C.2y-6
D.2y+6

Found 2 solutions by ankor@dixie-net.com, ewatrrr:
Answer by ankor@dixie-net.com(22740)

Ali will be 6y years old in 6 years time.
6y - 6 = present age
How old was Ali 4y years ago?
6y - 4y - 6 = 2y - 6

Answer by ewatrrr(24785)

Let x be Ali's age NOW
x + 6 = 6y
x = 6y-6
......substitute for x
4y years ago...
x - 4y = 6y-6 - 4y = 2y -6